OpenStudy (anonymous):

Whoever is good at precalc pleaseee help!!!! find the domain of the function: f(x)=√2-x the square root is over the x too

OpenStudy (watchmath):

Since the square root only non-negative inputs, then we need \(2-x\geq 0\) i.e., \(x\leq 2\)

OpenStudy (watchmath):

only accepts non-negative inputs

OpenStudy (anonymous):

want to see the choices?

OpenStudy (watchmath):

In interval notation that is equicalent to \((-\infty,2]\).

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

for the given functions f(x)=8x=5, g(x)=4x-1, find the compostion (fog)(x) the o btw the f and g is supposed to be a little of i think meaning of